[Date Prev][Date Next][Thread Prev][Thread Next][Search] [Main Index] [Thread Index] [HEASARC Mailing List Archives]

CGRO Biweekly Status Report

  Compton Observatory Science Report #169, Monday October 28, 1994
      Chris Shrader, Compton Observatory Science Support Center

         Questions or comments can be sent to the CGRO SSC.
          Phone: 301/286-8434
          e-mail:   NSI_DECnet: GROSSC::SHRADER 
          Internet: shrader@grossc.gsfc.nasa.gov

                          Spacecraft Status

The Observatory, scientific instruments and all spacecraft subsystems
are currently functioning in a nearly flawless manner.

We note that the responsibilities of the Compton GRO Project
Operations Director were transferred from Robert Wilson to Keith
Walyus on October 17. Robert Wilson, who has served in that role
since launch, is retiring on November 3. Best wishes to Bob and many
thanks for his years of outstanding service to the project!

                         Instrument Reports

EGRET operations were normal this biweekly period. Delivery of data
to the CGRO-SSC remains on scheduled. Interactions with Guest
Investigators remains at a good level. 

A paper giving very detailed results on the Vela Pulsar appeared in
Astronomy and Astrophysics (Kanbach et al., A&A 289, 855). Energy
spectra as a function of phase in the period are provided allowing a
comparison to pulsar theories that was previously not possible. 

Several EGRET papers will be presented at the forthcoming HEAD
meeting, including an invited talk on the diffuse radiation and a
paper presenting the evolution of gamma-ray bright blazars based on
the gamma-ray data alone.


OSSE operations are normal.

In viewing period 402.0 (18-25 Oct), the Z-axis target was the
galactic plane near (l,b) = (310,0) (Guest Investigator J. Skibo),
and the X-axis target was 3C279 (PI team).

In viewing period 402.5 (25 Oct - 1 Nov), the Z-axis target is again
the galactic plane (l,b) = (310,0) (Guest Investigator J. Skibo), and
the X-axis target is LMC X-1 (PI team).

Several more interesting OSSE results have recently been accepted for
publication in ApJ Part 1 or 2 and will be distributed as preprints
shortly. L.-S. The et al. report that OSSE upper limits on Ti-44 line
emission from Cas A are somewhat lower than the flux reported by
COMPTEL. Grabelsky et al. give spectra from the black hole candidate
GX339-4 in outburst and quiescence, with a strong indication of a
direct correspondence between the low X-ray state and gamma-ray
outbursts. Smith et al. report very stringent upper limits on
Compton-back-scattered annihilation radiation from the galactic
center covering >180 days of observation.

Data from viewing period 228 were delivered to the Compton GRO
Science Support Center archive this week. The targets were SN1993J
and the Vela Pulsar.


The COMPTEL instrument is performing well and continues routine

The collaboration will be well represented at next week's HEAD
meeting in Napa Valley, California. In addition to an invited talk by
R. Diehl on gamma-ray line spectroscopy with GRO, there will be
presentations on active Galactic nuclei, x-ray binaries, and
gamma-ray bursts observed by COMPTEL.

During the week following the HEAD conference there will be a general
team meeting of the collaboration at MPE in Garching, Germany.


Many of the BATSE team scientists will be presenting papers at the
HEAD meeting next week at Napa. Most are currently working hard to
finish preparations for their presentations. 

The x-ray transients GRO J1719-24 = GRS 1716-249 and GX 339-4 are
still in outburst.

The following sources were detected by the BATSE pulsed source
monitor in the past two weeks: Cen X-3, 4U 1626-67, 2S 1417-624, GX
1+4, Vela X-1, and GX 301-2. The outburst of 2S 1417-624 has now
lasted 63 days.

As of October 27th, BATSE has detected 1150 cosmic gamma-ray bursts
out of a total of 3155 on-board triggers in 1283 days of operation.
There have been 740 triggers due to solar flares with emission above
60 keV. Since the BATSE burst trigger criteria were changed, there
have been four triggers due to atmospheric gamma-ray flashes.